OpenStudy (nightrider):

4) Ori drove 203.75 miles while on vacation. Assuming he only drove at the set speeds 55 miles per hour and 40 miles per hour, how long did he travel at each speed if he drove for 4.25 hours? (Hint: Make a table.) Answer: Ori drove 2.25 hours at 55 mph and 2 hours at 40 mph I need to know how to get to that answer using a table, and I'm really struggling.

OpenStudy (welshfella):

u can solve it using a system of equations

OpenStudy (welshfella):

let x = distance travelled at 55 mph and y = distance at 40 mph and t1 be the time at 55 and t2 the time at 40 mph we have x + y = 207.33 and x = 55t1 and y = 40t2 so substituting;- 55t1 + 40t2 = 207.33 ----------------(1) also:- t1 + t2 = 4.25 .......................................(2) Sole equations (1) and (2) to find the required times.

OpenStudy (mhchen):

2016-09-24T18:22:51+00:00 X=number of hours, when Ori drove at 55 mph y=number of hours, when Ori drove at 40 mph We can suggest the next system of equations: x+y=4.25 55x+40y=203.75 We solve this system of equation by substitution method. x+y=4.25 ⇒y=(4.25-x) 55x+40(4.25-x)=203.75 55x+170-40x=203.75 15x=203.75-170 15x=33.75 x=33.75/15=2.25 y=(4.25-x)=4.25-2.25=2

OpenStudy (mhchen):

Now you see how it's solved? Think of what the answer will be after the next step.

OpenStudy (mhchen):

You gave me a medal :(